Miller and Sons' static budget for 9,600 units of production includes $37,400 for direct materials, $45,000 for direct labor, variable utilities of $6,000, and supervisor salaries of $16,300. A flexible budget for 12,000 units of production would show

Round your final answer to the nearest dollar. Do not round interim calculations.

a.direct materials of $46,750, direct labor of $56,250, utilities of $7,500, and supervisor salaries of $19,560

b.total variable costs of $104,700

c.direct materials of $46,750, direct labor of $56,250, utilities of $7,500, and supervisor salaries of $16,300

d.the same cost structure in total
Answer #1

Direct materials - (37,400/9,600) * 12,000 = 46,750

Dirscf labor - (45,000/9,600) * 12,000 = 56,250

Utilitues - (6,000/9,600) * 12,000 = 7,500

Supervisor salary (Fixed) - 16,300

Option C